Home‎ > ‎Crew 165‎ > ‎

0309-DailySummaryReport

Hello Mission Support!

Daily Summary Report (DSR) for 09 March 2016 and our Sol 3 on Mars!

SOL: 3
Person filling out Report: Ken Sullivan
Summary Title: Busy interior work
Mission Status: All systems nominal

Sol 2 Activity Summary: Food measurement research continues. A lot of small activities, NorCal Rover batteries, cables, toilet checks, photo/video work, Thesis work, cleaning, organizing, HAL, and Japanese food!

Anomalies in work: None 

Weather: Mostly sunny, light wind.  

Look ahead plan: EVA 3, ATV repair, soil samples 

Crew Physical Status: Good!

EVA: N/A 

Reports to be filed:
- Commander Report
- Engineering Report
- Journalist Report
- Daily Summary Report
- EVA #3 Request
- HSO Report

Support Requested: We need water badly. Down to 8.5 inches, not sure what the unusable water level is.  No ATV fuel in jerry cans. We also need Teflon tape.  Engineering has advised to check filter but we have no teflon tape nor did the prior crew.